40 students scored 70 on a test the standard deviation is 5.5 if the test of 10 students are picked randomly what is the probability their mean is greater than 77
& what is their probabilty their mean is less than 77

Answer by stanbon(75887)

40 students scored 70 on a test the standard deviation is 5.5 if the test of 10 students are picked randomly what is the probability their mean is greater than 77
& what is their probabilty their mean is less than 77
------
Using population size = 40 and pop. std = 5.5,
pop mean = np = 70
----
For all samples of size 10, the mean of the sample means is 70
and the std is 5.5/sqrt(10)
-----
Problem:
z(77) = (77-70)/(5.5/sqrt(10)) = 1.7393
p(x-bar > 77) = normalcdf(1.7303,100) = 0.0410
===================
p(x-bar < 77) = 1 - 0.0410 = 0.9590
===========================================
Cheers,
Stan H.